# On a conjecture concerning the Fisher–Widom line and the line of vanishing excess isothermal compressibility in simple fluids

### v1

@@ 4 Discussion
In this paper we have addressed one aspect of the role played by the attractive and repulsive parts of the intermolecular potential on the thermodynamic and structural properties of fluids. In particular, we have dealt with a conjecture, introduced by Stopper et al. [17], concerning the proximity of the FW line and the line of vanishing excess isothermal susceptibility (for which we have coined the name Seno line) in simple fluids. To test the validity of such a conjecture, we have taken the intermolecular potential to be the Jagla potential [25], since this model potential may account for multiple fluid transitions and for some of the thermodynamic and dynamic anomalies observed in water. Both the one-dimensional and the three-dimensional fluids have been considered. The second virial coefficient and the Zeno line, which also reflect the role played by the attractive and repulsive parts of the potential, have been obtained for these model fluids too. For the sake of illustration, we have taken in the two systems the set of parameters displayed in Equation (6).
The consideration of the one-dimensional system allowed us to derive exact results for all three lines. In this instance, we find that the conjecture is not satisfied (cf. Figure 1). Since the Seno line is defined by the condition $\int_{0}^{\infty}\mathrm{d}r\,r^{n}h(r)=0$, with $n=0$ for one-dimensional systems, one might reasonably argue that a modified condition with $n>0$ would emphasise the attractive part of the interaction and could serve as a better proxy for the FW line. However, our findings (not shown) indicate that $n=1$ and $n=2$ produce just the opposite effect.
In the case of the three-dimensional system, we have obtained approximate theoretical results with the RFA approach and we have also carried out MC simulations. Our findings indicate that, in contrast to what we found for the one-dimensional Jagla fluid, the conjecture of Ref. [17] is satisfied reasonably well, at least for $\rho^{*}=0.20$, $0.25$, $0.30$, $0.35$ and $0.4$. Interestingly, in the same range we also find a proximity between all three lines. Whether this feature will hold also for other fluids is worth investigating. On the other hand, we also find that, while the RFA approach agrees qualitatively in the description of the density behaviour of the FW, Seno and Zeno lines, it overestimates in general the points on the Zeno line and fails to capture the proximity of the Zeno line with the other two lines in the density interval mentioned above. Furthermore, our analysis confirms that although the RFA approach provides generally good results for the structural and thermodynamic quantities, it exhibits poor performance in accurately predicting the behaviour of the three transition lines, especially under conditions of high density and/or low temperature.
@@ Acknowledgments
A.M.M., S.B.Y. and A.S. acknowledge financial support from Grant No. PID2020-112936GB-I00 funded by MCIN/AEI/10.13039/501100011033, and from Grant No. IB20079 funded by Junta de Extremadura (Spain) and by ERDF “A way of making Europe.” A.M.M. is also grateful to the Spanish Ministerio de Ciencia e Innovación for a predoctoral fellowship PRE2021-097702. A.R.R. acknowledges financial support from Grant No. PID2021-126348NB-I00 funded by MCIN/AEI/10.13039/501100011033. M.L.H. thanks Universidad de Extremadura for its hospitality during his sabbatical year. The authors are grateful to the computing facilities of the Instituto de Computación Científica Avanzada of the University of Extremadura (ICCAEx) and of C3UPO (University Pablo de Olavide), where the simulations were run.
@@
Appendix A Low-temperature limit of the Zeno, Seno and FW lines for the one-dimensional fluid
In this Appendix we consider the one-dimensional Jagla fluid and analyze the limit $\beta^{*}\to\infty$ of the Zeno, Seno and FW lines, proving that all of them end at the Boyle density $\rho_{B}=\lambda_{1}^{-1}$.
@@ A.2 Seno line
Finally, Equation (8) gives $Z\to\lambda_{1}\beta p$, i.e., $\rho\to\lambda_{1}^{-1}$.
@@ A.3 FW line
In this case we have to deal with Equations (11). Taking the limit $\beta^{*}\to\infty$, one can see that $\beta p\to 0$ and $\kappa-\beta p\to 0$. Then, taking into account Equation (22), Equation (11c) yields

### v2

@@ 4 Discussion
In this paper we have addressed one aspect of the role played by the attractive and repulsive parts of the intermolecular potential on the thermodynamic and structural properties of fluids. In particular, we have dealt with a conjecture, introduced by Stopper et al. [17], concerning the proximity of the FW line and the line of vanishing excess isothermal compressibility (for which we have coined the name Seno line) in simple fluids. To test the validity of such a conjecture, we have taken the intermolecular potential to be the Jagla potential [26], since this model potential may account for multiple fluid transitions and for some of the thermodynamic and dynamic anomalies observed in water. Both the one-dimensional and the three-dimensional fluids have been considered. The second virial coefficient and the Zeno line, which also reflect the role played by the attractive and repulsive parts of the potential, have been obtained for these model fluids too. For the sake of illustration, we have taken in the two systems the set of parameters displayed in Equation (6).
The consideration of the one-dimensional system allowed us to derive exact results for all four lines. In this instance, we find that the conjecture is not satisfied (cf. Figure 1). Since the Seno line is defined by the condition $\int_{0}^{\infty}\mathrm{d}r\,r^{n}h(r)=0$, with $n=0$ for one-dimensional systems, one might reasonably wonder whether a modified condition with $n>0$ would emphasise the attractive part of the interaction and could serve as a better proxy for the FW line. However, our findings (not shown) indicate that $n=1$ and $n=2$ produce just the opposite effect.
In the case of the three-dimensional system, we have obtained approximate theoretical results with the RFA approach and we have also carried out MC simulations. Our findings indicate that, in contrast to what we found for the one-dimensional Jagla fluid, the conjecture of Ref. [17] is satisfied reasonably well, at least for $\rho^{*}=0.20$, $0.25$, $0.30$, $0.35$ and $0.4$. This is not very surprising, since the criterion of using the ideal-gas-like isothermal compressibility to estimate the FW line [17] is actually a mean-field idea that should work best when higher dimensions are considered. Interestingly, in the same density range we also find a proximity between the Zeno, Seno and FW lines. Whether this feature will hold also for other fluids is worth investigating. On the other hand, we also find that, while the RFA approach agrees qualitatively in the description of the density behaviour of the FW, Seno and Zeno lines, it overestimates in general the points on the Zeno line and fails to capture the proximity of the Zeno line with the other two lines in the density interval mentioned above. Furthermore, our analysis confirms that, although the RFA approach provides generally good results for the structural and thermodynamic quantities, it exhibits poor performance in accurately predicting the behaviour of the three transition lines, especially under conditions of high density and/or low temperature.
@@ Acknowledgments
The authors are grateful to the computing facilities of the Instituto de Computación Científica Avanzada of the University of Extremadura (ICCAEx) and of C3UPO (University Pablo de Olavide), where the simulations were run. M.L.H. thanks Universidad de Extremadura for its hospitality during his sabbatical year.
@@
Funding
A.M.M., S.B.Y., A.S. and M.L.H. acknowledge financial support from Grant No. PID2020-112936GB-I00 funded by MCIN/AEI/10.13039/501100011033, and from Grant No. IB20079 funded by Junta de Extremadura (Spain) and by ERDF “A way of making Europe.” A.M.M. is also grateful to the Spanish Ministerio de Ciencia e Innovación for a predoctoral fellowship PRE2021-097702. A.R.R. acknowledges financial support from Grant No. PID2021-126348NB-I00 funded by MCIN/AEI/10.13039/501100011033.
Appendix A Low-temperature limit of the Zeno, Seno, FW and Widom lines for the one-dimensional fluid
In this Appendix we consider the one-dimensional Jagla fluid and analyze the limit $\beta^{*}\to\infty$ of the Zeno, Seno, FW and Widom lines, proving that the first three of them end at the Boyle density $\rho_{B}=\lambda_{1}^{-1}$, while the Widom line ends at $\rho_{B}/2$.
@@ A.2 Seno line
Finally, Equation (8) gives $Z\to\lambda_{1}\beta p$, i.e. $\rho\to\lambda_{1}^{-1}$.
@@ A.3 FW line
In this case, we have to deal with Equations (11). Taking the limit $\beta^{*}\to\infty$, one can see that $\beta p\to 0$ and $\kappa-\beta p\to 0$. Then, taking into account Equation (23), Equation (11c) yields
@@
A.4 Widom line
In the low-temperature regime, the Widom line necessarily resides below the FW line, thereby rendering the damping coefficient $\kappa$ determined by Equation (11c). For a fixed value of $\beta p$, it can be seen that $\kappa-\beta p\sim\beta^{*}e^{-\beta^{*}}$ as $\beta^{*}\to\infty$. Consequently, by substituting $\Omega(-\kappa+\beta p)\to-(\kappa-\beta p)^{-1}+ae^{\beta^{*}}/\beta^{*}$ and $\Omega(\beta p)\to ae^{-\beta p\lambda_{1}}e^{\beta^{*}}/\beta^{*}$ into Equation (11c), we obtain
$\kappa\to\beta p+\frac{\beta^{*}e^{-\beta^{*}}}{a}\left(1-e^{-\beta p\lambda_{1}}\right)^{-1}.$ (30)
Now, the Widom condition $(\partial\kappa/\partial\beta p)_{\beta}=0$ yields
$\beta p\to\sqrt{\frac{\beta^{*}}{a\lambda_{1}}}e^{-\beta^{*}/2}.$ (31)
Therefore, $\Omega(\beta p)\to a{e^{\beta^{*}}}/{\beta^{*}}$, $\Omega^{\prime}(\beta p)\to-2a\lambda_{1}{e^{\beta^{*}}}/{\beta^{*}}$, which implies $\rho\to(2\lambda_{1})^{-1}$, i.e. half the Boyle density. This result is analogous to the one previously obtained for the one-dimensional triangle-well fluid [18].